OpenStudy (anonymous):

just help me set it up and I'll solve it :) Find the volume of the solid obtained by rotating the region bounded by the given curves about the specified axis. y=x^2, y=0, x=2, about the y-axis

OpenStudy (anonymous):

If you visualize the graph of y (where y=0 and x=2 are boundaries), you will see that its domain is in the first quadrant from y = 0 to 4. We need it in terms of y if we want to use the disc method, so x = sqrt{y}. Therefore integrate \[\pi \int\limits_{0}^{4}(\sqrt{y})dy\]

OpenStudy (anonymous):

I got 12pi?

OpenStudy (anonymous):

I had 16pi/3

OpenStudy (anonymous):

nope thats wrong... mine is too

OpenStudy (anonymous):

I forgot to square the radius. Integrate y.

OpenStudy (anonymous):

pi(y^2/2) from 0 to 4

OpenStudy (anonymous):

8pi